# Topological frequency conversion in rhombohedral multilayer graphene

### v1

@@ Appendix C Estimate of dissipation rate from nonzero chemical potential
Here we estimate the contribution to dissipation from a nonzero chemical potential $\mu$—that is, from the presence of an equilibrium Fermi surface. We consider the regime where the characteristic relaxation time is much larger than the oscillation period of the driving modes, $\omega\tau\gg 1$ and neglect the effects of the perpendicular displacement field on the band dispersion. We moreover assume $\mu$ to be small, working to leading order in $\mu$. Without loss of generality, we consider the case where $\mu>0$ (negative $\mu$ is related to our results below through particle hole symmetry).
Eq. (10) of the main text implies that the dissipation rate is given by
$\eta_{\rm diss}(t)=\frac{\langle E(t)\rangle-\langle E_{\rm eq}(t)\rangle}{\tau}$ (22)
where
$\langle E(t)\rangle$ $=$ $\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,{\rm Tr}[\hat{\rho}(\mathbf{k},t)\hat{H}(\mathbf{k},t)],$ (23)
$\langle E_{\rm eq}(t)\rangle$ $=$ $\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,{\rm Tr}[\hat{\rho}_{\rm eq}(\mathbf{k},t)\hat{H}(\mathbf{k},t)].$ (24)
Here $\hat{\cdot}$ indicate operators acting on second-quantized Bloch space.
@@
Since $\hat{H}(\mathbf{k},t)=\hat{H}(\mathbf{k}+e{\bf{A}}(t)/\hbar)$, we have $\hat{\rho}_{\rm eq}(\mathbf{k},t)=\hat{\rho}_{\rm eq}(\mathbf{k}+e{\bf{A}}(t)/\hbar)$. Thus, a shift of integration variable in Eq. (24), $\mathbf{k}\to\mathbf{k}+e{\bf{A}}(t)/\hbar$ shows that $\langle E_{\rm eq}(t)\rangle$ is time-independent and given by
$E_{\rm eq}=\sum_{n}\int\!\!\frac{d^{2}\mathbf{k}^{\prime}}{4\pi^{2}}f_{\rm\beta}(\varepsilon_{n}(\mathbf{k}^{\prime})-\mu)\varepsilon_{n}(\mathbf{k}^{\prime}).$ (25)
@@
Assuming adiabatic dynamics, and working in the limit $\omega\tau\gg 1$, $\rho(\mathbf{k},t)$ is diagonal in the eigenbasis of $H$, with stationary eigenvalues $p_{n}(\mathbf{k})=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}dtf_{\beta}(\varepsilon_{n}(\mathbf{k}+e{\bf{A}}(t)/\hbar)$ [31]. As a result,
$\langle E(t)\rangle=\sum_{n}\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,p_{n}(\mathbf{k})\varepsilon_{n}(\mathbf{k}+e{\bf{A}}(t)/\hbar)$ (26)
@@
$\bar{\eta}_{\rm mr}=\frac{1}{\tau}(\bar{E}-E_{\rm eq}),$ (27)
with $\bar{E}\equiv\lim_{T\to\infty}\int_{0}^{T}\frac{dt}{T}\langle E(t)\rangle$ denoting the average total energy of all electrons in the system. We compute this quantity by inserting the expression for $p_{n}(\mathbf{k})$ in the definition for $\bar{E}$. Using a change integration variables, $\mathbf{k}\to\mathbf{k}-e{\bf{A}}(t)/\hbar$, we obtain
$\bar{E}=\sum_{n}\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,\bar{p}_{i}(\mathbf{k})\varepsilon_{n}(\mathbf{k})$ (28)
@@
$\bar{p}_{n}(\mathbf{k})=\lim_{T_{i}\to\infty}\int_{0}^{T_{1}}\frac{dt_{1}}{T_{1}}\int_{0}^{T_{2}}\frac{dt_{2}}{T_{2}}f_{\beta}\left(\varepsilon_{n}\left(\mathbf{k}+\frac{e}{\hbar}[{\bf{A}}(t_{1})-{\bf{A}}(t_{2})]\right)-\mu\right)$ (29)
@@
$\bar{\eta}_{\rm mr}=\frac{1}{\tau}\sum_{n}\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,\varepsilon_{n}(\mathbf{k})[\bar{p}_{n}(\mathbf{k})-f_{\beta}(\varepsilon_{n}-\mu)]$ (30)
To estimate $\bar{\eta}_{\rm mr}$, we approximate $\bar{p}_{0}(\mathbf{k})=1$, and assume the valence band carriers to be uniformly smeared over a disk of radius $\Delta k$ where $\Delta k=3eA_{\parallel}/2\hbar$ with $A_{\parallel}$ denoting the characteristic scale of the in-plane vector potential from the drive (recall that we consider the limit of small $\mu$, where $k_{\rm F}$ in equilibrium is much smaller than $\Delta k$). This results in $\bar{p}_{1}(\mathbf{k})=\frac{k_{\rm F}^{2}}{\Delta k^{2}}\theta(\Delta k-|\mathbf{k}|)$ with $\theta$ denoting the step function. The estimate of $\bar{p}_{1}(\mathbf{k})$ is based on the fact that $\bar{p}(\mathbf{k})$ has maximum at $|\mathbf{k}|=0$, decreases monotonically with $|\mathbf{k}|$, and vanishes for $|\mathbf{k}|\geq 2eA_{\parallel}/\hbar$. We finally assume $f_{\beta}(\varepsilon-\mu)\propto\theta(-(\varepsilon-\mu))$.
@@
$\bar{\eta}_{\rm mr}\approx\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,\frac{\gamma_{L}}{\tau}\mathbf{k}^{L}\left[\theta(\Delta-|\mathbf{k}|)\frac{k_{\rm F}^{2}}{\Delta k^{2}}-\theta(k_{\rm F}-|\mathbf{k}|)\right]$ (31)
@@
$\bar{\eta}_{\rm mr}\approx\frac{\mu k_{\rm F}^{2}}{2\pi\tau(L+2)}[(\Delta k/k_{\rm F})^{L}-1]$ (33)
@@
$\bar{\eta}_{\rm mr}\approx\mu^{2/L}\frac{\gamma_{L}^{2/L-1}}{2\pi\tau(L+2)}\left(\frac{3eA_{\parallel}}{\hbar}\right)^{L}$ (35)
@@
$\bar{\eta}_{\rm mr}\approx\frac{\mu^{2/L}\gamma_{L}^{1-2/L}}{2\pi\tau(L+2)}\left(\frac{3eE\cos\theta}{\hbar\omega}\right)^{L}x$ (36)
@@
$\bar{\eta}_{\rm mr}\approx\frac{\mu}{8\pi\tau}\left(\frac{3eE\cos\theta}{\hbar\omega}\right)^{2}\quad(L=2),$ (37)
We now estimate the dissipation caused by the Fermi surface from this expression, and compare with numerical simulations carried out at non-zero $\mu$ in Figs. 7, 8. Note that Fig. 8 shows the total dissipated power, including both Landau-Zener absorption and momentum relaxation due to the presence of a Fermi surface. We expect that dissipation strength at $\mu=0$ serves as a good proxy for the characteristic rate of Landau-Zener absorption, while the difference in dissipation to the $\mu=0$ value, ${\cal I}_{\rm diss}^{+}(\mu)-{\cal I}_{\rm diss}^{+}(0)$ serves as a reasonable estimate for the momentum-relaxation induced dissipation, $\eta_{\rm mr}$.
[…]
This analysis also predicts that Fermi-surface-induced dissipation stays constant when scaling up the frequencies in conjunction with the incidence angle such that $\cos\theta_{j}/\omega_{j}$ remains fixed, as in Fig. 3 of the main text. We also verify this scaling in Fig. 8. While the $\mu=0$ dissipation increases with frequency due to more prominent Landau-Zener tunneling, the Fermi surface contribution (which turns on with $\mu$) is on the same scale and roughly independent of frequency. Note that we consider incommensurate frequencies; for commensurate frequencies, we expect the dissipation from Landau-Zener absorption to be further suppressed as demonstrated in Fig. 3.
Figure 7: Scaling of the (time-averaged) frequency conversion power ${\cal P}^{+}_{\rm FC}$ (panels a–d), and the corresponding dissipated power ${\cal P}^{+}_{\rm diss}$ (panels e–h), as a function of the chemical potential $\mu$. Color bars are expressed in units of the power quantum $hf_{1}f_{2}$. We consider incommensurate frequencies with ratio $r=1/\sqrt{2}$ and parameters identical to those used for Fig. 2a–c in the main text. As the chemical potential is increased, two trends are observed. First, the quantization of the frequency conversion power is gradually lost due to the opposite contribution of the conduction band (panels a–d). Second, the dissipation profile acquires a diffuse contribution from the underlying Fermi surface (panels e–h), which coexists with the sharp “ring of fire” due to Landau-Zener absorption.
@@
Figure 8: Scaling of the valley-resolved frequency conversion intensity ${\cal I}^{+}_{\rm FC}$ (left) and the corresponding dissipated power ${\cal I}^{+}_{\rm diss}$ (right), as a function of the chemical potential $\mu$ and different frequencies $\Omega=\sqrt{\omega_{1}\omega_{2}}$ as in Figs. 2 and 7. The data for the lowest $\Omega$ are obtained by integrating the power distributions shown in Fig. 7 (a-d) and (e-h), respectively.
Appendix D Estimate of the heating rate from driving
[…]

### v2

@@ Appendix D Estimate of dissipation rate with a non-zero chemical potential
Here we estimate the contribution to dissipation from a state with a non-zero chemical potential $\mu$—that is, from the presence of an equilibrium Fermi surface. We consider the regime where the characteristic relaxation time is much larger than the oscillation period of the driving modes, $\omega_{j}\tau\gg 1$ and neglect the effects of the perpendicular displacement field on the band dispersion. We moreover assume $\mu$ to be small, working to leading order in $\mu$. Without loss of generality, we consider the case where $\mu>0$ (negative $\mu$ is related to our results below through the approximate particle hole symmetry of the band structure of RMG).
Eq. (9) of the main text implies that the dissipation intensity is given by
$\eta_{\rm diss}(t)=\frac{\langle\mathcal{E}(t)\rangle-\langle\mathcal{E}_{\rm eq}(t)\rangle}{\tau},$ (21)
with
$\langle\mathcal{E}(t)\rangle$ $=$ $\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,{\rm Tr}[\hat{\rho}(\mathbf{k},t)\hat{H}(\mathbf{k},t)],$ (22)
$\langle\mathcal{E}_{\rm eq}(t)\rangle$ $=$ $\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,{\rm Tr}[\hat{\rho}_{\rm eq}(\mathbf{k},t)\hat{H}(\mathbf{k},t)],$ (23)
denoting the instantaneous and equilibrium energy densities, respectively. Here $\hat{\cdot}$ indicate operators acting on the Fock space of fermions with a given wavevector $\mathbf{k}$.
@@
Since $\hat{H}(\mathbf{k},t)=\hat{H}(\mathbf{k}+e{\bf{A}}(t)/\hbar)$, we have $\hat{\rho}_{\rm eq}(\mathbf{k},t)=\hat{\rho}_{\rm eq}(\mathbf{k}+e{\bf{A}}(t)/\hbar)$. Thus, a shift of integration variable in Eq. (23), $\mathbf{k}\to\mathbf{k}+e{\bf{A}}(t)/\hbar$ shows that $\langle\mathcal{E}_{\rm eq}(t)\rangle$ is time-independent and given by
$\mathcal{E}_{\rm eq}=\sum_{n}\int\!\!\frac{d^{2}\mathbf{k}^{\prime}}{4\pi^{2}}f_{\rm\beta}(\varepsilon_{n}(\mathbf{k}^{\prime})-\mu)\varepsilon_{n}(\mathbf{k}^{\prime}).$ (24)
@@
Assuming adiabatic dynamics, and working in the limit $\omega_{j}\tau\gg 1$, $\rho(\mathbf{k},t)$ is diagonal in the eigenbasis of $H$, with stationary eigenvalues $p_{n}(\mathbf{k})=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}dtf_{\beta}(\varepsilon_{n}(\mathbf{k}+e{\bf{A}}(t)/\hbar)$ [31]. As a result,
$\langle\mathcal{E}(t)\rangle=\sum_{n}\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,p_{n}(\mathbf{k})\varepsilon_{n}(\mathbf{k}+e{\bf{A}}(t)/\hbar)$ (25)
@@
$\bar{\eta}_{\rm mr}=\frac{1}{\tau}(\bar{\mathcal{E}}-\mathcal{E}_{\rm eq}),$ (26)
with ¯$\bar{\mathcal{E}}\equiv\lim_{T\to\infty}\int_{0}^{T}\frac{dt}{T}\langle\mathcal{E}(t)\rangle$ denoting the average energy density of the system. We compute this quantity by inserting the expression for $p_{n}(\mathbf{k})$ in the definition for $\bar{\mathcal{E}}$. Using a change integration variables, $\mathbf{k}\to\mathbf{k}-e{\bf{A}}(t)/\hbar$, we obtain
¯$\bar{\mathcal{E}}=\sum_{n}\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,\bar{p}_{i}(\mathbf{k})\varepsilon_{n}(\mathbf{k}),$ (27)
@@
$\bar{p}_{n}(\mathbf{k})=\lim_{T_{i}\to\infty}\int_{0}^{T_{1}}\frac{dt_{1}}{T_{1}}\int_{0}^{T_{2}}\frac{dt_{2}}{T_{2}}f_{\beta}\left(\varepsilon_{n}\left(\mathbf{k}+\frac{e}{\hbar}[{\bf{A}}(t_{1})-{\bf{A}}(t_{2})]\right)-\mu\right).$ (28)
@@
$\bar{\eta}_{\rm mr}=\frac{1}{\tau}\sum_{n}\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,\varepsilon_{n}(\mathbf{k})[\bar{p}_{n}(\mathbf{k})-f_{\beta}(\varepsilon_{n}-\mu)].$ (29)
To estimate $\bar{\eta}_{\rm mr}$, we approximate $\bar{p}_{0}(\mathbf{k})=1$, and assume the valence band carriers to be uniformly smeared over a disk of radius $\Delta k=3eA_{\parallel}/2\hbar$ with $A_{\parallel}$ denoting the characteristic scale of the in-plane vector potential from the drive (recall that we consider the limit of small $\mu$, where $k_{\rm F}$ in equilibrium is much smaller than $\Delta k$). This results in $\bar{p}_{1}(\mathbf{k})=\frac{k_{\rm F}^{2}}{\Delta k^{2}}\theta(\Delta k-|\mathbf{k}|)$ with $\theta$ denoting the Heaviside step function. The estimate of $\bar{p}_{1}(\mathbf{k})$ is based on the fact that $\bar{p}(\mathbf{k})$ has maximum at $|\mathbf{k}|=0$, decreases monotonically with $|\mathbf{k}|$, and vanishes for $|\mathbf{k}|\geq 2eA_{\parallel}/\hbar$. We finally assume $f_{\beta}(\varepsilon-\mu)\propto\theta(-(\varepsilon-\mu))$.
Fig. S3: Scaling of the (time-averaged) frequency conversion power ${\cal P}^{+}_{\rm FC}$ (panels a–d), and the corresponding dissipated power ${\cal P}^{+}_{\rm diss}$ (panels e–h), as a function of the chemical potential $\mu$. Color bars are expressed in units of the power quantum $hf_{1}f_{2}$. We consider incommensurate frequencies with ratio $r=1/\sqrt{2}$ and parameters identical to those used for Fig. 2a–c in the main text. As the chemical potential is increased, two trends are observed. First, the quantization of the frequency conversion power is gradually lost due to the opposite contribution of the conduction band (panels a–d). Second, the dissipation profile acquires a diffuse contribution from the underlying Fermi surface (panels e–h), which coexists with the sharp “ring of fire” due to Landau-Zener absorption.
@@
$\bar{\eta}_{\rm mr}\approx\int\!\frac{d^{2}\mathbf{k}}{4\pi^{2}}\,\frac{\gamma_{L}}{\tau}\mathbf{k}^{L}\left[\theta(\Delta-|\mathbf{k}|)\frac{k_{\rm F}^{2}}{\Delta k^{2}}-\theta(k_{\rm F}-|\mathbf{k}|)\right].$ (30)
@@
$\bar{\eta}_{\rm mr}\approx\frac{\mu k_{\rm F}^{2}}{2\pi\tau(L+2)}[(\Delta k/k_{\rm F})^{L}-1].$ (32)
@@
$\bar{\eta}_{\rm mr}\approx\mu^{2/L}\frac{\gamma_{L}^{2/L-1}}{2\pi\tau(L+2)}\left(\frac{3eA_{\parallel}}{\hbar}\right)^{L}.$ (34)
@@
$\bar{\eta}_{\rm mr}\approx\frac{\mu^{2/L}\gamma_{L}^{1-2/L}}{2\pi\tau(L+2)}\left(\frac{3eE\cos\theta}{\hbar\omega}\right)^{L}x.$ (35)
@@
$\bar{\eta}_{\rm mr}\approx\frac{\mu}{8\pi\tau}\left(\frac{3eE\cos\theta}{\hbar\omega}\right)^{2}\quad(L=2).$ (36)
@@
Fig. S4: Scaling of the valley-resolved frequency conversion intensity ${\cal I}^{+}_{\rm FC}$ (left) and the corresponding dissipation intensity ${\cal I}^{+}_{\rm diss}$ (right), as a function of the chemical potential $\mu$ and different frequencies $\Omega=\sqrt{\omega_{1}\omega_{2}}$, with an incommensurate ratio $r=1/\sqrt{2}$ as in Figs. 2, S1 and S3. The data for the lowest $\Omega$ are obtained by integrating the power distributions shown in Fig. S3 (a-d) and (e-h), respectively.
We now estimate the dissipation caused by the Fermi surface from this expression, and compare with numerical simulations carried out at non-zero $\mu$ in Figs. S3, S4. Note that Fig. S4 shows the total dissipated power, including both Landau-Zener absorption and momentum relaxation due to the presence of a Fermi surface. We expect that dissipation strength at $\mu=0$ serves as a good proxy for the characteristic rate of Landau-Zener absorption, while the difference in dissipation to the $\mu=0$ value, ${\cal I}_{\rm diss}^{+}(\mu)-{\cal I}_{\rm diss}^{+}(0)$ serves as a reasonable estimate for the momentum-relaxation induced dissipation, $\eta_{\rm mr}$.
Using frequencies $\omega\sim 0.68$ THz, electric fields $E\sim 0.1$ V/nm, relaxation time $\tau=500$ ps, incidence angles $\theta_{j}-\pi/2\approx 0.1^{\circ}$, our estimate in Eq. (36) yields $\eta_{\rm mr}\approx 0.5$ W/mm2 for $\mu\approx 40$ meV. This rough estimate agrees reasonably well with the numerical simulations shown in Fig. S4, which suggest $\eta_{\rm mr}\sim 2$ W/mm2.
This analysis also predicts that Fermi-surface-induced dissipation stays constant when scaling up the frequencies in conjunction with the incidence angle such that $\cos\theta_{j}/\omega_{j}$ remains fixed, as in Fig. 3 of the main text. We also verify this scaling in Fig. S4. While the $\mu=0$ dissipation increases with frequency due to more prominent Landau-Zener tunneling, the Fermi surface contribution (which turns on with $\mu$) is on the same scale and roughly independent of frequency. Note that we consider incommensurate frequencies; for commensurate frequencies, we expect the dissipation from Landau-Zener absorption to be further suppressed as demonstrated in Fig. 3.
Appendix E Estimate of the heating rate from driving
We finally estimate the heating rate from the driving configuration we consider in the main text, which will limit the lifetime of topological frequency conversion. From the numerical simulations in Fig. 3 we extract a typical scale for the dissipation intensity $I_{\rm diss}\sim 40$ W/mm2 for commensurate driving with frequencies $f_{i}\sim 1$ THz and $\tau=500$ ps, accounting for dissipation from both valleys. The heat capacity of graphene in the temperature range $T\sim 10-100$K is dominated by phonons and of order $C\sim\left(0.1-1\right)\frac{\rm J}{\rm K\,mol}$ [71]. Using the density of carbon atoms of $2/A_{\rm uc}$ with the unit cell area $A_{\rm u.c.}=5.2\,{\rm\AA}^{2}$ yields the heat capacity per unit area $c\sim 10^{-11}-10^{-12}\frac{\rm J}{\rm K\,mm^{2}}$. We use this value to estimate the heat capacity for bilayer graphene as $2c$, resulting in a heating rate given by $\eta=\frac{I_{\rm diss}}{2c}\sim 10^{11}-10^{12}$ K/s. The phonon temperature in irradiated bilayer graphene is thus expected to increase by a Kelvin for every $1-10$ ps. This estimate should allow for multiple THz-frequency cycles before the temperature increase of phonons causes the electronic lifetime $\tau$ to decrease substantially—thus resulting in the gradual disappearance of frequency conversion.